Anti-SPAM rules on mail server modified to allow VPN clients to relay mail through the University mail server.
The /home volume on the mail server filled from 97% to 99% in about 6 hours. In order to prevent problems with the mail server, data was moved from the /home volume to /home3. The LDAP directory was updated and permissions were reassigned to the relocated data. The /home volume is currently at 75% capacity.
The time on houdini, primary DNS server, drifted 10 minutes out, causing a failure in synchronizing with the secondary DNS server, harry. Time was reset. Josh will install NTP synchronization.
The time in the Library Windows domain drifted off this morning. Tardis 2000 was applied to keep the time synchronized.
Adjustments to the PIX firewall rules were made to allow the ezproxy server to function, and to allow encoding and management of the streaming media server.
Found an improper MX record in the main zone in the external DNS.
The Micros (cash register) server was restarted